OpenStudy (anonymous):

can someone remind me how to do these: 125=5^x

OpenStudy (anonymous):

wolframalpha.com walks you through it

OpenStudy (anonymous):

ok, what are they even called? I didn't know how to search for it

OpenStudy (anonymous):

just put your equation in

OpenStudy (anonymous):

I'll just teach you...

OpenStudy (anonymous):

You have to take the log or natural log (ln) of both sides

OpenStudy (anonymous):

\[\log 125 = \log 5^x\]

OpenStudy (anonymous):

This is equal to \[\log 125 = x \log 5\]

OpenStudy (anonymous):

I think it's just common sense. 5x5=25 25x5 = 125 therefore 5x5x5=125 therefore 5^3 = 125 therefore X=3

OpenStudy (anonymous):

the exponent can be brought out in front.

OpenStudy (nubeer):

@Anna1998 x is in the power not multiply with 5

OpenStudy (anonymous):

@ChmE when I do log of 125 I get 2.0969

OpenStudy (anonymous):

@brandonloves not finished yet

OpenStudy (anonymous):

shutting up :-)

OpenStudy (anonymous):

now to get x by itself we divide both sides by log 5\[\frac{ \log 125 }{ \log 5 }=x\]

OpenStudy (anonymous):

3!

OpenStudy (anonymous):

\[125=5^x\]\[\log 125 = \log 5^x\]\[\log 125 = x \log 5\]\[\frac{ \log 125 }{ \log 5 }=x\]

OpenStudy (anonymous):

seems like I remember doing them another way that was pretty easy... what if it was 16^x=28 .83?

OpenStudy (anonymous):

oh, maybe we just guessed them...

OpenStudy (anonymous):

 this is how we worked them in class, but that was a couple months ago, and I forgot what it all means @ChmE